OpenStudy (anonymous):

is x=1 a solution of the equation 2-8x=-6

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

2-8*1 =-6 = RHS = -6 so it is a solution

OpenStudy (anonymous):

2 + 6 = 8x there fore x =1 is a solution

OpenStudy (anonymous):

When you all get finished please look at mine

OpenStudy (anonymous):

2-8x = -6 => -8x = -8 => x = 1

OpenStudy (anonymous):

yes